Find three consecutive odd numbers whose sum is 147 ?
Answer by lenny460(1073)

Find three consecutive odd numbers whose sum is 147 ?
Le the three consecutive odd numbers be:
n + 1
n + 3 and
n + 5
Therefore:
n + 1 + n + 3 + n + 5 = 147
Combine like terms:
n + n + n = 3n
1 + 3 + 5 = 9
3n + 9 = 147
Subtract 9 from each side of the equation.
3n + 9 - 9 = 147 - 9
9 - 9 = 0
147 - 9 = 138
3n = 138
n = 46
Therefore:
n + 1 = 46 + 1 = 47
n + 3 = 46 + 3 = 49
n + 5 = 46 + 5 = 51

Answer:
The three consecutive odd numbers are:
47, 49 and 51













Lennox Obuong
Algebra Tutor
Nairobi, Kenya
Email: lennoxobuong@yahoo.com